OpenStudy (anonymous):

Find the remaining trigonometric ratios sec (phi)= −2.5, pi/2

OpenStudy (anonymous):



OpenStudy (anonymous):

there is an angle whose secant is \(\frac{5}{2}=2.5\) ignoring the signs for a moment, find the third side by pythagoras it is \[\sqrt{5^2-2^2}=\sqrt{25-4}=\sqrt{21}\] then use the appropriate ratios, making sure to keep track of the signs you should have in quadrant 2

OpenStudy (anonymous):

thank youu :)